# Understanding and Supporting Formal Email Exchange by Answering AI-Generated Questions

### v1

@@ Abstract.
Replying to formal emails is time-consuming and cognitively demanding, as it requires polite phrasing and ensuring an adequate response to the sender’s demands. Although systems with Large Language Models (LLM) were designed to simplify the email replying process, users still needed to provide detailed prompts to obtain the expected output. Therefore, we proposed and evaluated an LLM-powered question-and-answer (QA)-based approach for users to reply to emails by answering a set of simple and short questions generated from the incoming email. We developed a prototype system, ResQ, and conducted controlled and field experiments with 12 and 8 participants. Our results demonstrated that QA-based approach improves the efficiency of replying to emails and reduces workload while maintaining email quality compared to a conventional prompt-based approach that requires users to craft appropriate prompts to obtain email drafts. We discuss how QA-based approach influences the email reply process and interpersonal relationship dynamics, as well as the opportunities and challenges associated with using a QA-based approach in AI-mediated communication.
@@
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714016isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
Email is a common tool for users to share information (Nelson et al., 2011; Whittaker et al., 2006) and manage tasks (Bellotti et al., 2005). It has been found that users spend an average of 28% of their workweek reading and replying to emails (Institute, 2012). However, for many, checking and responding to emails is time-consuming and cognitively demanding (Mark et al., 2016). Responding to emails, especially in cultures that value courteous email exchanges, requires users to understand the sender’s requests and compose polite messages that reflect the sender’s intentions. This involves considering various elements such as tone, style, diction, and structure (Chen, 2015). Putting effort into constructing courteous emails and responding promptly is important because the lack of these elements can result in a negative perception by recipients (Resendes et al., 2012; Vignovic and Thompson, 2010; Kalman and Rafaeli, 2011), potentially harming trust and damaging relationships in relationships in formal communication settings. In this paper, we define formal email exchange as a type of communication where the email structure is typically clear and organized, employing polite and respectful language with a specific focus. Examples of formal email exchange include office-related communication, research collaboration in academic institutions, and interactions with external organizations.
@@
To address this issue, we propose a QA-based approach in which the system analyzes incoming emails and generates questions that invite users to respond to efficiently create the desired draft. In this paper, we define and implement a QA-based approach by generating questions based on the text of the email. Then, the system creates an email draft based on the users’ answers to these questions. This QA-based approach was motivated by prior studies suggesting that answering structured questions helps users articulate their needs more effectively (Kim et al., 2024; Jannach et al., 2022; Cao et al., 2023). This approach aims to lower the cognitive burden of drafting replies, help users quickly understand the sender’s requests, and reduce the need for cumbersome prompting by breaking down the prompt-generation process into smaller, more manageable QA steps.
@@ 2.3. The Psychological and Relational Impact of AIMC Tools
Previous research has extensively discussed the impact of AIMC tools on the psychological aspects of both recipients and senders, as well as their relationships. These effects can be broadly categorized into two areas: (1) The impact on the recipient and sender relationships (2) The impact on the sender’s self-perception.
@@ 4. Proposed LLM-Powered QA-Based Approach: ResQ
Fig. 2 illustrates the overview of the process of a reply message using ResQ. Fig. 3 shows the actual interface of ResQ. The following sections describe the specific functions involved in each step of this process.
@@
When the first activates ResQ to reply to an email, the system uses an LLM (in this study, GPT-4o (OpenAI, 2024b)) to generate multiple-choice questions (Fig. 3-C). Additionally, when users click on the generated question, the relevant part of the email is highlighted (Fig. 3-A). The LLM then extracts all parts that require a user’s reply, presents the possible response options, and generates a question for the user. Following the approach described in (Bsharat et al., 2024), we designed a structured prompt that guided the LLM to determine the number of questions necessary and sufficient to cover all the requirements of the incoming email without omission. Instead of pre-specifying a fixed number of questions, the prompt instructed the model to produce an “appropriate” number of questions, where “appropriate” is defined as the minimal set of questions needed to address all points raised by the sender while avoiding redundant or irrelevant inquiries. To ensure that the questions were generated systematically rather than randomly, we provided explicit criteria within the prompt. These criteria included referencing the sender’s intent, quoting relevant portions of the original email verbatim, and offering multiple-choice options where applicable. We also provided concrete examples within the prompt to illustrate the desired format and style of the generated questions and corresponding answer choices. By doing so, we ensured that the LLM’s output was both well-grounded and easy for the recipient to respond to. The detailed prompt used to guide the LLM in this process is shown in the appendix.
@@ 5.1. Experiment Design
Study 1 was designed to quantitatively assess how ResQ influences the writing process (RQ1), quality of email replies (RQ2), and perceived relationships with others (RQ3) compared to scenarios without AI intervention and when using traditional AIMC tools. The experiment targeted Japanese participants and was conducted entirely in Japanese. This experiment context was designed as communication in formal settings, such as office-related communication, research collaboration in academic institutions, and interactions with external organizations. It focused on time-consuming emails that were characterized as lengthy, containing multiple requests, or requiring detailed and polite responses. Simple and straightforward emails, such as those that can be answered with a single word or phrase (e.g., “Understood”), were excluded from the scope.
Participants were assigned the role of message recipients and were tasked with crafting replies based on the scenarios and supplementary information provided. The messages used in the experiment were collected from 10 volunteers, who provided actual emails they had received in formal communication contexts. These volunteers included office workers, graduate students, and teaching staff, all of whom were Japanese and engaged in email communication regularly (at least once per month). To ensure anonymity, identifying details were removed during the preparation process. Based on the design principles of ResQ, we excluded extremely short emails and emails that could be replied to with a single word from the selection process. Each scenario included details about the sender, the recipient, and the context in which the message was received. Multiple scenarios were included in the experiment to minimize the influence of any single scenario and increase the variety. Additionally, supplementary information, such as the recipient’s schedule and potential questions, was provided to prevent excessive variability in the responses among participants.
@@ 5.4. Procedure
The participants first read the study instructions and the right to participate and then consented to participate in the experiment. Next, they were given an explanation of the purpose of the experiment and the use of the AI systems (Prompt-based and QA-based systems). Participants were then randomly assigned to reply to six emails per condition using a Latin square design, which counterbalanced the order of conditions and mitigated potential order effects  [^2] We conducted analyses to examine the potential order effect. The results of this analysis are provided in the appendix.. In each condition, participants first engaged in a practice session where they read and replied to two emails to familiarize themselves with the system. Then, they read and replied to six emails, which were presented in a randomly assigned order to further reduce any sequence-related biases. After replying to six emails for each condition, participants were asked to complete a questionnaire regarding their experience with the task. To ensure participants could manage their workload during the study, they were allowed to take a short break after completing tasks for each condition. After completing all conditions, they were asked to fill out a comparative questionnaire evaluating the three conditions. In addition, follow-up interviews were conducted to gather deeper insights into their experiences and preferences. This study was conducted remotely for all participants and lasted approximately two and a half hours in total.
@@ 5.5. Evaluation Session
After completing the main experiment, we conducted an additional evaluation session to assess the quality of the email responses created by participants and the impressions of participants as email senders. This session involved a group of eighteen Japanese evaluators (ten males and eight females, aged 20-57) recruited via a local participant recruiting platform  [^3] Participants were recruited from Lancers.jp, an online freelancing platform.. The average age of the evaluators was 40.6 (SD = 8.3). The evaluators had a minimum of five years and an average of 18.7 years of experience in email-based communication. Furthermore, with the exception of one individual, the evaluators engaged in email-based communication at least once a month. Each evaluator assessed reply emails written by twelve different participants for a specific scenario. The evaluators were paid approximately \$2.5 USD for their participation, and the evaluation session lasted around fifteen minutes.
@@ 6.1.1. Efficiency of Replying to Emails (H1-a)
First, we compared the efficiency of replying to emails across three conditions. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measure ANOVA showed that there was a significant difference in participants’ efficiency of replying to emails across three conditions ($F[2,22]=14.8$, $p<0.001$, $\eta_{p}^{2}=0.57$). Post-hoc analysis with Holm correction revealed that participants’ efficiency of replying to emails in the QA-based condition was significantly higher compared to both the No-AI $(t(11),p=0.002{\color[rgb]{0,0,0},d=1.38})$ and the Prompt-based $(t(11),p=0.046{\color[rgb]{0,0,0},d=0.65})$ conditions. Thus, H1-a was supported. The QA-based approach enhanced participants’ email replying efficiency.
@@ 6.1.3. Cognitive Load for Replying to Emails (H1-b)
The results of the Raw-TLX are shown in Fig. 5. According to the one-way repeated measure ANOVA with Greenhouse-Geisser correction, there was a significant difference in participants’ cognitive load for replying to emails among the three conditions $(F[1.1,12.1]=12.6,p=0.003{\color[rgb]{0,0,0},\eta_{p}^{2}=0.53})$. Post-hoc analysis with Holm correction revealed that participants’ cognitive load for replying to emails in the QA-based condition was significantly lower compared to both the No-AI $(t(11),p=0.008{\color[rgb]{0,0,0},d=1.12})$ and Prompt-based $(t(11),p=0.018{\color[rgb]{0,0,0},d=0.81})$ conditions. Therefore, H1-b was supported. The QA-based approach reduced participants’ cognitive workload while replying to the emails.
@@ 6.1.4. Difficulty in Understanding Email Content (H1-b)
] Additionally, H1-b was also supported by the questionnaire survey results (Fig. 6 H1-b). The Friedman test revealed a significant difference among the three conditions in terms of understanding the sender’s intent and requests $(\chi^{2}(2)=10.6,p=0.005{\color[rgb]{0,0,0},W=0.44})$. Post-hoc analysis using the Durbin-Conover test with Holm correction showed that participants in the QA-based condition found it significantly easier to understand the sender’s intent and requests compared to those in both No-AI $(p=0.003{\color[rgb]{0,0,0},r=0.61})$ and Prompt-based $(p=0.005{\color[rgb]{0,0,0},r=0.73})$ conditions.
@@ 6.1.5. Satisfaction with Completing Task (H1-c)
The results of the satisfaction with completing participants’ tasks are shown in Fig. 6, H1-c. The two items measuring satisfaction showed high internal consistency, with a Cronbach’s Alpha of $0.889$. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measure ANOVA showed that there was a significant difference in participants’ satisfaction with completing tasks across three conditions $(F[2,22],p<0.001{\color[rgb]{0,0,0},\eta_{p}^{2}=0.79})$. Post-hoc analysis with Holm correction revealed that participants’ satisfaction with completing tasks in the QA-based condition was significantly higher compared to both the No-AI ($t(11)$, $p<0.001$, $d=2.39$) and the Prompt-based ($t(11)$, $p=0.029$, $d=0.72$) conditions. Therefore, H1-c was supported. The QA-based approach improved participants’ satisfaction with completing their tasks while replying to the emails.
@@ 6.1.6. Difficulty in Initiating the Action for Replying to Emails (H1-d)
The questionnaire survey results about participants’ difficulty in initiating the action for replying to emails are shown in Fig. 6, in H1-d. According to the Friedman test, a significant difference in participants’ difficulty in initiating the action for replying to emails was observed among the three conditions $(\chi^{2}(2)=19.8,p<0.001{\color[rgb]{0,0,0},W=0.83})$. Post-hoc analysis using the Durbin-Conover test with Holm correction revealed that participants in the QA-based condition perceived significantly higher barriers to initiating email response tasks than those in the No-AI $(p<0.001{\color[rgb]{0,0,0},r=0.85})$ and Prompt-based $(p<0.001{\color[rgb]{0,0,0},r=0.68})$ conditions. Therefore, H1-d was supported. The QA-based approach reduced participants’ difficulties in initiating the action to reply to emails.
@@ 6.1.7. Sense of Agency and Control (H1-e)
The questionnaire survey results about a sense of agency and control are shown in Fig. 6, H1-e. The Friedman test revealed a significant difference among the three conditions for both the sense of agency $(\chi^{2}(2)=22.8,p<0.001{\color[rgb]{0,0,0},W=0.95})$ and the sense of control $(\chi^{2}(2)=21.3,p<0.001$,
${\color[rgb]{0,0,0}W=0.89})$. Post-hoc analysis using the Durbin-Conover test with Holm correction showed that participants in the QA-based condition found it significantly reduced participants’ sense of agency compared to both the No-AI $(p<0.001{\color[rgb]{0,0,0},r=0.88})$ and the Prompt-based $(p<0.001{\color[rgb]{0,0,0},r=0.77})$ conditions. Additionally, post-hoc analysis using the Durbin-Conover test with Holm correction showed that participants in the QA-based condition found it significantly reduced participants’ sense of control compared to both the No-AI $(p<0.001{\color[rgb]{0,0,0},r=0.88})$ and the Prompt-based $(p=0.006{\color[rgb]{0,0,0},r=0.56})$ conditions. Thus, H1-e was supported. The QA-based approach reduced participants’ sense of agency and sense of control while replying to the emails.

### v2

@@ Abstract.
Replying to formal emails is time-consuming and cognitively demanding, as it requires crafting polite phrasing and providing an adequate response to the sender’s demands. Although systems with Large Language Models (LLMs) were designed to simplify the email replying process, users still need to provide detailed prompts to obtain the expected output. Therefore, we proposed and evaluated an LLM-powered question-and-answer (QA)-based approach for users to reply to emails by answering a set of simple and short questions generated from the incoming email. We developed a prototype system, ResQ, and conducted controlled and field experiments with 12 and 8 participants. Our results demonstrated that the QA-based approach improves the efficiency of replying to emails and reduces workload while maintaining email quality, compared to a conventional prompt-based approach that requires users to craft appropriate prompts to obtain email drafts. We discuss how the QA-based approach influences the email reply process and interpersonal relationship dynamics, as well as the opportunities and challenges associated with using a QA-based approach in AI-mediated communication.
@@
copyright: acmlicensedjournalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714016isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
Email is a common tool for users to share information (Nelson et al., 2011; Whittaker et al., 2006) and manage tasks (Bellotti et al., 2005). It has been found that users spend an average of 28% of their workweek reading and replying to emails (Institute, 2012). However, for many, checking and responding to emails is time-consuming and cognitively demanding (Mark et al., 2016). Responding to emails, especially in cultures that value courteous email exchanges, requires users to understand the sender’s requests and compose polite messages that reflect the sender’s intentions. This involves considering various elements such as tone, style, diction, and structure (Chen, 2015). Putting effort into constructing courteous emails and responding promptly is important because the lack of these elements can result in a negative perception by recipients (Resendes et al., 2012; Vignovic and Thompson, 2010; Kalman and Rafaeli, 2011), potentially harming trust and damaging relationships in formal communication settings. In this paper, we define formal email exchange as a type of communication where the email structure is typically clear and organized, employing polite and respectful language with a specific focus. Examples of formal email exchange include office-related communication, research collaboration in academic institutions, and interactions with external organizations.
@@
To address this issue, we propose a QA-based approach in which the system analyzes incoming emails and generates questions that invite users to respond to efficiently create the desired draft. In this paper, we define and implement a QA-based approach by generating questions based on the text of the email. Then, the system creates an email draft based on the users’ answers to these questions. This QA-based approach was motivated by prior studies suggesting that answering structured questions helps users articulate their needs more effectively (Kim et al., 2024; Jannach et al., 2022; Cao et al., 2023). This approach aims to lower the cognitive burden of drafting replies, help users quickly understand the sender’s requests, and reduce the need of cumbersome prompting by breaking down the prompt-generation process into smaller, more manageable QA steps.
@@ 2.3. The Psychological and Relational Impact of AIMC Tools
Previous research has extensively discussed the impact of AIMC tools on the psychological aspects of both recipients and senders, as well as on their relationships. These effects can be broadly categorized into two areas: (1) The impact on the recipient and sender relationships (2) The impact on the sender’s self-perception.
@@ 4. Proposed LLM-Powered QA-Based Approach: ResQ
Fig. 2 illustrates the overview of how a reply message is created using ResQ. Fig. 3 shows the actual interface of ResQ. The following sections describe the specific functions involved in each step of this process.
@@
When a user first activates ResQ, the system uses an LLM (in this study, GPT-4o (OpenAI, 2024b)) to generate multiple-choice questions (Fig. 3-C). The LLM extracts all parts of the email that require a reply, generates corresponding questions and presents possible response options. Additionally, if a user clicks on any generated question, the relevant part of the email is highlighted (Fig. 3-A).
Following the approach described in (Bsharat et al., 2024), we designed a structured prompt that guides the LLM to determine how many questions are necessary and sufficient to cover all requirements in the incoming email without omission. Instead of pre-specifying a fixed number of questions, the prompt instructs the model to produce an “appropriate” number of questions, where “appropriate” is defined as the minimal set of questions needed to address all points raised by the sender while avoiding redundant or irrelevant inquiries. To ensure that the questions were generated systematically rather than randomly, we provided explicit criteria within the prompt. These criteria included referencing the sender’s intent, quoting relevant portions of the original email verbatim, and offering multiple-choice options where applicable. We also provided concrete examples within the prompt to illustrate the desired format and style of the generated questions and corresponding answer choices. By doing so, we ensured that the LLM’s output was both well-grounded and easy for the recipient to answer. The detailed prompt used to guide the LLM in this process is shown in the appendix.
@@ 5.1. Experiment Design
Study 1 was designed to quantitatively assess how ResQ influences the writing process (RQ1), the quality of email replies (RQ2), and the perceived relationships with others (RQ3) compared to scenarios without AI intervention and when using traditional AIMC tools. The experiment targeted Japanese participants and was conducted entirely in Japanese. This experimental context was designed as communication in formal settings, such as office-related communication, research collaboration in academic institutions, and interactions with external organizations. It focused on time-consuming emails that were characterized as lengthy, containing multiple requests, or requiring detailed and polite responses. Simple and straightforward emails, such as those that can be answered with a single word or phrase (e.g., “Understood”), were excluded from the scope.
Participants were assigned the role of message recipients and required to craft replies based on the scenarios and supplementary information provided. The messages used in the experiment were collected from 10 volunteers who provided real emails they had received in formal communication contexts. These volunteers included office workers, graduate students, and teaching staff, all of whom were Japanese and engaged in email communication regularly (at least once per month). To ensure anonymity, identifying details were removed during the preparation process. Based on the design principles of ResQ, we excluded extremely short emails and emails that could be replied to with a single word from the selection process. Each scenario included details about the sender, the recipient, and the context in which the message was received. Multiple scenarios were included in the experiment to minimize the influence of any single scenario and increase the variety. Additionally, supplementary information, such as the recipient’s schedule and potential questions, was provided to prevent excessive variability in the responses among participants.
@@ 5.4. Procedure
The participants first read the study instructions and the right to participate and then consented to participate in the experiment. Next, they were given an explanation of the purpose of the experiment and the use of the AI systems (Prompt-based and QA-based systems). Participants were then randomly assigned to reply to six emails per condition using a Latin square design, which counterbalanced the order of conditions and mitigated potential order effects  [^2] We conducted analyses to examine the potential order effect. The results of this analysis are provided in the appendix.. In each condition, participants first engaged in a practice session where they read and replied to two emails to familiarize themselves with the system. Then, they read and replied to six emails, which were presented in a randomly assigned order to further reduce any sequence-related biases. After replying to six emails for each condition, participants were asked to complete a questionnaire regarding their experience with the task. To ensure participants could manage their workload during the study, they were allowed to take a short break after completing tasks in each condition. After completing all conditions, they were asked to fill out a comparative questionnaire evaluating the three conditions. In addition, follow-up interviews were conducted to gather deeper insights into their experiences and preferences. This study was conducted remotely for all participants and lasted approximately two and a half hours in total.
@@ 5.5. Evaluation Session
After completing the main experiment, we conducted an additional evaluation session to assess the quality of the email responses created by participants and the impressions of participants as email senders. This session involved a group of eighteen Japanese evaluators (ten males and eight females, aged 20-57) recruited via a local participant recruiting platform  [^3] Participants were recruited from Lancers.jp, an online freelancing platform.. The average age of the evaluators was 40.6 (SD = 8.3). The evaluators had a minimum of five years and an average of 18.7 years of experience in email-based communication. Furthermore, with the exception of one individual, the evaluators engaged in email-based communication at least once a month. Each evaluator assessed email replies written by twelve different participants for a specific scenario. The evaluators were paid approximately \$2.5 USD for their participation, and the evaluation session lasted around fifteen minutes.
@@ 6.1.1. Efficiency of Replying to Emails (H1-a)
First, we compared the efficiency of replying to emails across three conditions. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measures ANOVA showed that there was a significant difference in participants’ efficiency of replying to emails across three conditions ($F[2,22]=14.8$, $p<0.001$, $\eta_{p}^{2}=0.57$). Post-hoc analysis with Holm correction revealed that participants’ efficiency of replying to emails in the QA-based condition was significantly higher compared to both the No-AI $(t(11),p=0.002{\color[rgb]{0,0,0},d=1.38})$ and the Prompt-based $(t(11),p=0.046{\color[rgb]{0,0,0},d=0.65})$ conditions. Thus, H1-a was supported. The QA-based approach enhanced participants’ email replying efficiency.
@@ 6.1.3. Cognitive Load for Replying to Emails (H1-b)
The results of the Raw-TLX are shown in Fig. 5. According to the one-way repeated measures ANOVA with Greenhouse-Geisser correction, there was a significant difference in participants’ cognitive load for replying to emails among the three conditions $(F[1.1,12.1]=12.6,p=0.003{\color[rgb]{0,0,0},\eta_{p}^{2}=0.53})$. Post-hoc analysis with Holm correction revealed that participants’ cognitive load for replying to emails in the QA-based condition was significantly lower compared to both the No-AI $(t(11),p=0.008{\color[rgb]{0,0,0},d=1.12})$ and Prompt-based $(t(11),p=0.018{\color[rgb]{0,0,0},d=0.81})$ conditions. Therefore, H1-b was supported. The QA-based approach reduced participants’ cognitive workload while replying to the emails.
@@ 6.1.4. Difficulty in Understanding Email Content (H1-b)
Additionally, H1-b was also supported by the questionnaire survey results (Fig. 6 H1-b). The Friedman test revealed a significant difference among the three conditions in terms of understanding the sender’s intent and requests $(\chi^{2}(2)=10.6,p=0.005{\color[rgb]{0,0,0},W=0.44})$. Post-hoc analysis using the Durbin-Conover test with Holm correction showed that participants in the QA-based condition found it significantly easier to understand the sender’s intent and requests compared to those in both No-AI $(p=0.003{\color[rgb]{0,0,0},r=0.61})$ and Prompt-based $(p=0.005{\color[rgb]{0,0,0},r=0.73})$ conditions.
@@ 6.1.5. Satisfaction with Completing Task (H1-c)
The results of the satisfaction with completing participants’ tasks are shown in Fig. 6, H1-c. The two items measuring satisfaction showed high internal consistency, with a Cronbach’s Alpha of $0.889$. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measures ANOVA showed that there was a significant difference in participants’ satisfaction with completing tasks across three conditions $(F[2,22],p<0.001{\color[rgb]{0,0,0},\eta_{p}^{2}=0.79})$. Post-hoc analysis with Holm correction revealed that participants’ satisfaction with completing tasks in the QA-based condition was significantly higher compared to both the No-AI ($t(11)$, $p<0.001$, $d=2.39$) and the Prompt-based ($t(11)$, $p=0.029$, $d=0.72$) conditions. Therefore, H1-c was supported. The QA-based approach improved participants’ satisfaction with completing their tasks while replying to the emails.
@@ 6.1.6. Difficulty in Initiating the Action for Replying to Emails (H1-d)
The questionnaire survey results about participants’ difficulty in initiating the action for replying to emails are shown in Fig. 6, in H1-d. According to the Friedman test, a significant difference in participants’ difficulty in initiating the action for replying to emails was observed among the three conditions $(\chi^{2}(2)=19.8,p<0.001{\color[rgb]{0,0,0},W=0.83})$. Post-hoc analysis using the Durbin-Conover test with Holm correction revealed that participants in the QA-based condition perceived significantly higher barriers to initiating email response tasks than those in the No-AI $(p<0.001{\color[rgb]{0,0,0},r=0.85})$ and Prompt-based $(p<0.001{\color[rgb]{0,0,0},r=0.68})$ conditions. Therefore, H1-d was supported. The QA-based approach reduced participants’ difficulty in initiating the action to reply to emails.
@@ 6.1.7. Sense of Agency and Control (H1-e)
The questionnaire survey results about a sense of agency and control are shown in Fig. 6, H1-e. The Friedman test revealed a significant difference among the three conditions for both the sense of agency $(\chi^{2}(2)=22.8,p<0.001{\color[rgb]{0,0,0},W=0.95})$ and the sense of control $(\chi^{2}(2)=21.3,p<0.001$, ${\color[rgb]{0,0,0}W=0.89})$. Post-hoc analysis using the Durbin-Conover test with Holm correction showed that participants in the QA-based condition found that it significantly reduced their sense of agency compared to both the No-AI $(p<0.001{\color[rgb]{0,0,0},r=0.88})$ and the Prompt-based $(p<0.001{\color[rgb]{0,0,0},r=0.77})$ conditions. Additionally, post-hoc analysis using the Durbin-Conover test with Holm correction showed that participants in the QA-based condition experienced a significantly reduction in their sense of control compared to both the No-AI $(p<0.001{\color[rgb]{0,0,0},r=0.88})$ and the Prompt-based $(p=0.006{\color[rgb]{0,0,0},r=0.56})$ conditions. Thus, H1-e was supported. The QA-based approach reduced participants’ sense of agency and sense of control while replying to the emails.